OpenStudy (anonymous):

Find the point of equilibrium for Demand Function: p = 37 − 0.0002x Supply Function: p = 22 + 0.00001x

OpenStudy (anonymous):

equilibrium means supply and demand are equal to each other

OpenStudy (anonymous):

okay @jayz657

OpenStudy (anonymous):

so set both of the equation equal to each other and solve for x that will be your answer

OpenStudy (anonymous):

1. (22.7143, 71428.6) 2. (71428.6, 71428.6) 3. (71428.6, 22.7143) 4. (22.7143, 22.7143) 5. None of these

OpenStudy (anonymous):

these are my choices

OpenStudy (anonymous):

@jayz657

OpenStudy (anonymous):

so i'm not sure

OpenStudy (anonymous):

show me your work first and i let you know if theres anything wrong first

OpenStudy (anonymous):

37 − 0.0002x= 22 + 0.00001x

OpenStudy (anonymous):

37-22= 15

OpenStudy (anonymous):

0.0002+0.00001= 0.00021

OpenStudy (anonymous):

71428.5714

OpenStudy (anonymous):

that's all I have

OpenStudy (anonymous):

okay so you have the value for x you just need to solve for p now

OpenStudy (anonymous):

all you need to to is plug into one of the equations and solve for p

OpenStudy (anonymous):

do*

OpenStudy (anonymous):

so plug your answer into the first equation to get your second value

OpenStudy (anonymous):

22.71428?

OpenStudy (anonymous):

yup thats rigth

OpenStudy (anonymous):

your answeer should be in this form btw (x,p)

OpenStudy (anonymous):

thank you so much!!!!

OpenStudy (anonymous):

can you help me with something else?

OpenStudy (anonymous):

yea sure post it here

OpenStudy (anonymous):

its on the pdf file number 20

OpenStudy (anonymous):

i dont see a number 20?

OpenStudy (anonymous):

nvm i found it give me a min to figure it out

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

from the picture you can do this find the area of the rectangle and subtract all the white right traigles to get the area of the shaded

OpenStudy (anonymous):